two point charges are kept separated by 4 cm of air and 6cm of a dielectric of relative permittivity 4. the equivalent dielectric separation between them so far their coulombian interaction is concerned is


a) 10cm  b) 8 cm  c) 5cm d) 16cm

Grade:12

2 Answers

Arun Kumar IIT Delhi
9 years ago
In a dielectric medium you can assume that the length in air is increased by


Thanks & Regards
Arun Kumar
IIT Delhi
Askiitians Faculty
Rafel
6 years ago
F in air will be equal to force in dielectric . (Kqq`)/r²=( Kqq`)/kx²............ k ->dielectric constantr= √k.xr=4 , k=4 x=2 Hence d=6 +2 = 8 cm